OpenStudy (anonymous):

Two numbers are in the ratio 5:2. One half of their sum is 10.5. Find the numbers.

OpenStudy (anonymous):

need help the answers are 15& 6 but dont know how to solve the problem ALGEBRA 1 PROBLREM

OpenStudy (sasogeek):

let the sum of the numbers = x (5/7)x + (2/7)x = 2(10.5) (5/7)x + (2/7)x = 21 solve for x and find the numbers (5/7)x and (2/7)x

OpenStudy (anonymous):

how

OpenStudy (anonymous):

15 and 6

OpenStudy (anonymous):

but how did u solve it

OpenStudy (anonymous):

since half of the sum is 10.5, the sum is 21. (10.5x2) so 5:2 is ratio of numbers, and u add 5+2 and it equals 7 so 21divided 7 equals 3and first number is 5x3 which i s15 and second number is 2x3=6